### v1

# Reasoning Efficiently Through Adaptive Chain-of-Thought Compression: A Self-Optimizing Framework

Abstract.
The introduction of Chain-of-Thought (CoT) reasoning has significantly enhanced the problem-solving capabilities of Large Language Models (LLMs) by prompting them to articulate intermediate reasoning steps rather than producing answers directly. This structured approach has led to significant improvements in accuracy and robustness across benchmarks in arithmetic, logic, and commonsense reasoning. Although effective, it introduces computational overhead. Each reasoning step requires additional token generation, which raises inference latency, memory usage, and key-value cache demands. These issues are particularly critical in domains such as software engineering, where deterministic logic and concise outputs are preferred.
To understand how CoT length impacts both model performance and user experience, we conduct an empirical study on widely used code generation datasets. We find that longer CoT does not always improve performance. Across benchmarks, excessively long reasoning often results in truncation, reduced accuracy, and latency increases of up to five times, with failed outputs consistently longer than successful ones. These findings challenge the assumption that longer reasoning is inherently beneficial and underscore the need for adaptive CoT control. Motivated by these findings, we propose SEER (Self-Enhancing Efficient Reasoning), an adaptive and self-enhancing framework for CoT compression that improves the conciseness of reasoning while preserving output quality, without relying on external compression tools. SEER follows an adaptive self-enhancing paradigm, where CoT responses are refined by combining Best-of-N sampling with task-aware adaptive filtering. The adaptive filter dynamically adjusts thresholds based on the token length of pre-inference model outputs, effectively controlling verbosity while reducing computational costs and preserving accuracy. We evaluate SEER on three software engineering tasks and one mathematical reasoning task. On average, SEER reduces CoT length by 42.1% without compromising performance, and accuracy even improves due to reduced truncation. Moreover, it substantially lowers truncation rates and eliminates most infinite reasoning loops, thereby enhancing both efficiency and robustness. These results highlight SEER as a practical approach to making the use of CoT more efficient, enabling CoT-enhanced LLMs to operate effectively even in resource-constrained settings and on complex tasks.
Chain-of-Thought (CoT), Large Language Models (LLMs), CoT Compression, Software Engineering, Reasoning Efficiency
copyright: none
1. Introduction
The introduction of Chain-of-Thought (CoT) marks a pivotal advancement in enhancing the reasoning capabilities of Large Language Models (LLMs), particularly for complex, multi-step tasks. By encouraging models to explicitly articulate intermediate reasoning steps, CoT enables a systematic, step-by-step breakdown of problems rather than merely producing an answer in a single attempt (Kojima et al., 2022; Wei et al., 2022). While CoT has demonstrated notable improvements in accuracy and robustness across tasks involving arithmetic, logic, and commonsense reasoning (Nye et al., 2021; Wang et al., 2023; Zhou et al., 2022), its effectiveness becomes even more pronounced in software engineering tasks. Such tasks typically demand precise logical reasoning, structured multi-step problem solving, and transparent decision-making, all of which are well supported by CoT. Recent work shows that CoT can substantially improve performance in multiple software engineering tasks (e.g., code generation, program synthesis, and automated debugging) by improving the interpretability of the model’s reasoning and aligning it with human problem-solving processes (Yao et al., 2023; Li et al., 2025).
However, this benefit comes with substantial costs. As most contemporary LLMs are based on the Transformer architecture (Vaswani et al., 2017), the attention mechanism incurs quadratic computational complexity with respect to context length, while the KV cache grows linearly with context size. Consequently, the verbose decoding of CoT significantly increases inference latency and memory usage, with KV-cache requirements emerging as a key bottleneck (Agrawal et al., 2024; Ho et al., 2024). This inefficiency is particularly critical in software engineering tasks, where timely responses are essential for developer productivity and effective decision-making. Moreover, standard CoT often “overthinks” even simple code generation tasks, generating excessive reasoning tokens despite the existence of straightforward solutions. In more complex scenarios (e.g., repairing software vulnerabilities), this overthinking can impair performance as it increases the inference costs and latencies (Cuadron et al., 2025). Consequently, reducing the length of CoT reasoning without compromising model performance remains a key challenge.
Recent research indicates that compressing the CoT in training data and subsequently fine-tuning the model can effectively manage the length of the generated output while still maintaining accuracy (Kang et al., 2025; Munkhbat et al., 2025a; Xia et al., 2025). Methods such as TokenSkip allow LLMs to eliminate semantically redundant tokens, achieving significant reductions (e.g., from 313 to 181 tokens on the GSM-8K dataset) with minimal performance loss (Xia et al., 2025). Advanced frameworks like Adaptive GoGI-Skip, R1-Compress, and CoLaR have further improved these gains by reducing CoT lengths by 45%–80% while preserving accuracy, as well as delivering 1.6 to 2 times speedups in inference (Zhuang et al., 2025; Wang et al., 2025b; Tan et al., 2025).
However, most of these methods are evaluated on mathematical reasoning datasets and commonsense QA, without assessing their effectiveness on software engineering problems; consequently, they may achieve good performance on mathematical reasoning tasks but exhibit substantially weaker results in software engineering contexts. Moreover, most of these approaches use external compression tools to compress CoT, such as GPT-4 in C3oT (Kang et al., 2025) and LLMLingua in TokenSkip (Xia et al., 2025), which may lead to information loss. Such information loss may lead to “thought leap” and ultimately influence the performance of the model (Xu et al., 2025b). For software engineering tasks, these compression tools may change the logic of the code or even make the code unable to execute.
In this work, we conduct an empirical study to investigate the impact of excessively long CoT outputs on coding tasks using four open-source reasoning models. Our empirical study shows that excessively long CoT reasoning can harm the performance of LLMs. On HumanEval, small LLMs suffer substantial accuracy drops from truncation (e.g., 60.36% to 46.34% on the 1.5B model), while larger models show limited gains since failed outputs are often linked to longer reasoning. In addition, lengthy CoT raises latency by up to five times, underscoring the need to control reasoning length. Motivated by these findings, we propose SEER (Self-Enhancing Efficient Reasoning), an adaptive and self-enhancing framework for CoT compression that improves the conciseness of reasoning while preserving output quality, without relying on external compression tools.
SEER adopts a self-enhancing paradigm, in which the model generates its own CoT-augmented responses and refines them via a two-stage data refinement process: 1) Best-of-N (BoN) sampling, which selects the shortest correct reasoning trace among multiple candidates to ensure both correctness and conciseness, and 2) an adaptive CoT filtering mechanism that dynamically calibrates maximum reasoning length based on the statistical properties of each dataset. This design enables task-specific control over verbosity, mitigating over-compression risks while eliminating pathological repetition patterns. Notably, SEER fine-tuned for one domain (e.g., code generation) also demonstrates strong transferability, preserving concise reasoning in other domains without retraining.
We evaluate SEER on three diverse software engineering tasks and one mathematical reasoning benchmark under an 8K token context constraint. Experimental results show that SEER achieves an average CoT length reduction of 42.1% while preserving or improving accuracy across all tasks. Compared to baselines such as TokenSkip and Naive BoN, our method remains robust in complex software engineering tasks, where other approaches often fail due to instability or logical corruption. Moreover, SEER mitigates infinite reasoning loops, which occur when a model repeatedly generates the same fragment until the context limit is reached, by up to 97.7%, substantially reducing truncation rates. In summary, our main contributions are as follows:
  • We conduct an empirical study on HumanEval to examine how excessively long CoTs affect coding tasks. The results show that overlong reasoning often causes truncation, loops, and latency, ultimately reducing accuracy and stability, motivating the need for effective CoT length control.
  • To address this, we propose SEER, a lightweight CoT compression framework that learns concise reasoning patterns from self-generated outputs, effectively reducing truncation and infinite loops without external modules.
  • Extensive experiments across software engineering and mathematical reasoning tasks demonstrate that SEER achieves an average 42.1% reduction in CoT length while maintaining or even improving pass@1 accuracy.
  • Ablation and cross-domain studies further confirm that adaptive CoT filtering is the key driver of efficiency and accuracy, with BoN sampling providing complementary benefits, and that SEER generalizes well across different task domains.
Paper Organization. The remainder of this paper is structured as follows: Section 2 reviews CoT efficiency challenges, the overthinking phenomenon, and self-generated fine-tuning. Section 3 presents a motivating empirical study on the impact of CoT length. Section 4 details the SEER framework, including BoN sampling and adaptive filtering. Section 5 outlines the datasets, baselines, and evaluation metrics. Section 6 reports experimental findings and ablation results, addressing performance, generalizability, and loop mitigation. Section 7 discusses broader implications and limitations, and Section 8 reviews related literature. Finally, Section 9 concludes the paper and outlines future research directions.
2. Background
2.1. Chain-of-Thought Efficiency
LLMs have achieved remarkable success in complex reasoning tasks, largely due to advanced prompting techniques like CoT (Wei et al., 2022). CoT elicits multi-step reasoning by instructing the model to generate a series of intermediate steps before producing a final answer. This explicit articulation of the thought process allows the model to deconstruct complex problems, track its progress, and correct intermediate errors, leading to significant performance gains in tasks that demand logical deduction, arithmetic calculation, and symbolic reasoning.
However, the verbosity of CoT introduces significant computational overhead. Each reasoning step requires additional token generation, directly increasing inference latency and resource consumption, particularly for the KV-cache (Agrawal et al., 2024; Ho et al., 2024). This overhead poses a critical bottleneck for real-time applications or systems with constrained computational budgets. Consequently, a key research direction in CoT efficiency is to strike an optimal balance: retaining the reasoning benefits of CoT while minimizing its associated costs. Approaches in this area typically explore methods such as pruning redundant steps, learning to generate more concise reasoning paths (Kang et al., 2025; Xia et al., 2025), or selectively applying CoT based on task complexity (Wang et al., 2025a).
2.2. Overthinking Phenomenon in CoT
Beyond simple inefficiency, the relationship between CoT length and accuracy is not a straightforward “longer is better.” Prior work shows that longer reasoning does not always translate into higher accuracy (Xie et al., 2025). Instead, many studies identify an “overthinking” phenomenon. Accuracy often improves with additional reasoning at the early stages, since the model has more opportunity to explore possible solution paths. However, once the reasoning exceeds an optimal range, performance begins to decline as the model produces redundant or even erroneous content (Chen et al., 2024b; Jin et al., 2024; Cuadron et al., 2025). This decline is a direct result of overthinking, where excessive reasoning introduces noise, deviates from the correct reasoning path, and ultimately undermines accuracy (Chen et al., 2024a). Taken together, these findings highlight that the key challenge is to carefully control CoT length so as to reduce overthinking while preserving accuracy.
2.3. Self-Generated Data for Self-Supervised Fine-Tuning
Recent research has demonstrated that LLMs can bootstrap their own capabilities by generating candidate data, applying automated quality filtering, and then fine-tuning on the filtered subset. STaR illustrates this approach. The model generates CoT rationales and answers, discards incorrect outputs, and fine-tunes iteratively on the correct rationale–answer pairs, reaching reasoning performance comparable to models that are thirty times larger (Zelikman et al., 2022). Additionally, extensions such as Quiet-STaR generalize STaR by generating token-level rationales during continued pretraining, improving zero-shot reasoning on benchmarks like GSM8K and CommonsenseQA without explicit task fine-tuning (Zelikman et al., 2024). These methods collectively form a cohesive pipeline: self-generation, automatic filtering, and self-fine-tuning. This structure allows LLMs to independently improve their reasoning, domain knowledge, and generalization capabilities without the need for human-labeled data.
3. Empirical Study
Before introducing our proposed framework, we conduct an empirical study to investigate how the length of CoT reasoning influences the performance, efficiency, and stability of LLMs in software engineering scenarios. Our analysis primarily targets coding tasks, a core application area of software engineering where precise and executable reasoning is essential. This study provides a quantitative assessment of the challenges introduced by excessively long CoT outputs, thereby motivating the design of our method.
3.1. Settings
3.1.1. Datasets
Our empirical study is conducted on the HumanEval (et al., 2021), which is a widely used benchmark consisting of manually curated Python programming problems. Each problem contains a function signature, a docstring, and a set of test cases for evaluating functional correctness. Additionally, in the final part of our empirical study, we use a more complex dataset, Codeforces (Penedo et al., 2025). Codeforces comprises challenging algorithmic optimization problems collected from one of the most popular websites among competitive programmers.
3.1.2. LLMs
We employ the DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025a) series of models to conduct our empirical study. This choice is motivated by the fact that it is one of the most popular reasoning models in recent times. Moreover, it is open-source, which facilitates deployment. We conduct experiments on four LLMs with sizes of 1.5B, 7B, 14B, and 32B.
3.1.3. Evaluation Metrics
To assess the performance of CoT, we employ two metrics.
[…]
Accuracy. We use Pass@1 as the primary metric to evaluate correctness. This metric calculates the proportion of problems for which the model produces a correct solution on the first attempt, directly reflecting its practical effectiveness in solving tasks.
[…]
Average Token Length. We report the average number of tokens generated in model outputs. This metric reflects the verbosity of the reasoning process and helps assess whether excessively long outputs introduce redundancy or cause truncation. It also facilitates a direct comparison of output lengths between models with and without CoT reasoning.
3.1.4. Implementation Details
[…]
3.2. Empirical Findings
[…]
Truncation Leads to Performance Degradation. Table 1 demonstrates that excessively long CoT outputs introduce severe truncation issues, which directly limit performance improvements. For the smallest model (DeepSeek-R1-Distill-Qwen-1.5B), accuracy drops sharply, with the pass@1 score falling from 60.36% to 46.34%. This loss coincides with an extreme average CoT length of 7,645 tokens. The underlying cause is frequent truncation, which removes critical parts of the solution. In fact, 63 out of 164 problems are truncated on this model. As the model size increases, the truncation problem becomes less frequent, and moderate gains emerge. The 7B and 14B models show performance gains with CoT, but these improvements are limited by long outputs and truncation. The 32B model still experiences truncation in about 8.5% of cases, which keeps its pass@1 unchanged at 90.24% with or without CoT.
Table 1. HumanEval Results of Reasoning Models (With vs. Without CoT)
Models HumanEval (With CoT) HumanEval (Without CoT)
Pass@1 (%) Avg Tokens Trunc Loop Pass@1 (%) Avg Tokens Trunc Loop
DeepSeek-R1-Distill-Qwen-1.5B 46.34 7,645 63 62 60.36 513 0 0
DeepSeek-R1-Distill-Qwen-7B 80.49 4,210 24 22 70.12 474 0 0
DeepSeek-R1-Distill-Qwen-14B 89.63 3,687 16 14 84.76 444 0 0
DeepSeek-R1-Distill-Qwen-32B 90.24 3,340 14 12 90.24 474 0 0
[…]
Finding 1:  Excessive CoT length induces harmful truncation, especially in small-scale LLMs. Small-scale LLMs are more easily affected by long CoTs because their context is quickly filled. For larger models, truncation is less frequent but still reduces the benefit, showing that longer reasoning does not always improve results once the context limit is reached.
[…]
Reasoning loop results in truncation. We manually examine truncation cases to better understand their impact on model performance. Our analysis shows that the model often falls into repetitive reasoning loops, repeatedly generating the same fragment. For example, it may continually output phrases such as “Wait, if …” until the maximum context length is reached. This behavior consumes a large number of tokens and severely reduces efficiency. As shown in Table 1, reasoning loops account for most truncation cases, with an average rate of 94.0%. When CoT generation is disabled, this problem disappears entirely. Larger models alleviate the issue to some extent, but reasoning loops are not fully eliminated.
These observations suggest that while CoT can enhance reasoning, uncontrolled generation often leads to loops and truncation, which harm both efficiency and reliability. Therefore, controlling CoT length is crucial to preserve its benefits without incurring instability.
[…]
Finding 2:  Reasoning loops are the dominant reason for truncation. They waste tokens, reduce efficiency, and persist across model scales, disappearing only when CoT is disabled. This highlights looping as a critical issue in CoT reasoning.
[…]
Is longer always better? While prior work (Jin et al., 2024) suggests that increasing the length of CoT reasoning improves accuracy, our findings reveal a different pattern. We conduct experiments on one of the most challenging problems in HumanEval (findmyway, 2024), namely HumanEval/129. Since the 1.5B model achieves almost no correct solutions in our experiments (pass@1 is nearly zero), we exclude it from further analysis. For each of the other models, we generate 10,000 responses and examine the distribution of CoT lengths, as shown in Figure 1, with the maximum token length set to 16,384. Contrary to the common assumption that longer reasoning consistently leads to better performance (Jin et al., 2024), our analysis reveals a clear pattern across all three models. Failed outputs are frequently associated with longer CoT lengths, while successful outputs are more concentrated around shorter lengths. For instance, in the 7B model, the median CoT length of failed generations (9,489 tokens) surpasses that of successful generations (8,296 tokens), a pattern that persists in both the 14B and 32B models. This suggests that excessively long reasoning chains may introduce redundancy, amplify error propagation, or even trigger context truncation, ultimately degrading performance. These results highlight that effective reasoning does not simply require “more thinking”, but rather appropriate reasoning lengths that balance sufficiency and conciseness.
Figure 1. Distribution of CoT lengths for passed and failed cases across DeepSeek-R1-Distill-Qwen models (7B, 14B, and 32B) on HumanEval/129. The results show that successful outputs do not necessarily require longer CoT reasoning. An excessively long reasoning chain can instead lead to a decline in accuracy.
[…]
Finding 3:  Longer CoT does not guarantee higher accuracy. Across all three models, failed generations consistently exhibit longer CoT lengths than successful ones (Figure 1). This indicates that excessively long reasoning chains often introduce redundancy and error propagation rather than improving performance, highlighting the need for appropriately controlled CoT lengths.
[…]
Long CoT introduces application latency. Excessively long CoT reasoning introduces significant latency overhead, which ultimately degrades user experience. To quantify this impact, we evaluate the inference performance of Deepseek-R1-Distill-Qwen models of varying sizes using the transformers library on a single NVIDIA A800 GPU. The measured generation speeds range from 46 tokens/s for the smallest (1.5B) model to 18 tokens/s for the largest (32B) model. Using these measurements, we estimate the average end-to-end latency on two benchmarks: HumanEval, representing relatively simple coding tasks, and Codeforces, which contains more complex algorithmic challenges. The results are presented in Figure 2.
Figure 2. Inference latency of different model sizes on HumanEval and Codeforces datasets with and without CoT reasoning. While CoT improves reasoning capability, it substantially increases application latency, particularly for larger models, posing practical constraints on real-world deployment.
Figure 2 shows that incorporating CoT substantially increases response time, particularly for larger models on complex tasks. On Codeforces, the 32B model with CoT requires an average of 776 seconds to complete the reasoning for each problem, compared to 157 seconds without CoT. Even the smaller 1.5B model takes more than five minutes per problem when using CoT. These delays make real-time use impractical. Moreover, all experiments were conducted under a 16k-token context limit, whereas Codeforces samples already exceed 14k tokens on average, implying that real-world latency would be even higher. Thus, while CoT improves reasoning, its verbosity imposes severe constraints on latency-sensitive applications, highlighting the need for effective CoT length control.
[…]
Finding 4:  Long CoT substantially slows down inference. For example, the 32B model on Codeforces requires nearly five times longer with CoT than without. This latency overhead makes CoT impractical for real-time or latency-sensitive applications, underscoring the need for effective length control.
4. Methodology
We propose SEER, a self-enhancing framework designed to enable LLMs to refine their reasoning capabilities by learning from their own generated outputs. Unlike conventional supervised fine-tuning pipelines that depend heavily on high-quality, human-labeled CoT annotations, SEER operates in a fully autonomous manner: the model generates candidate reasoning traces, filters them based on correctness and conciseness, and then learns from the selected high-quality examples. As illustrated in Figure 3, the framework is composed of three key stages: (1) Pre-inference generation of CoT responses, (2) BoN sampling for high-quality reasoning path selection, and (3) Adaptive CoT filtering to control verbosity without sacrificing reasoning fidelity. Through this design, SEER significantly reduces the reliance on prompt engineering and external compression tools, while promoting reasoning traces that are both accurate and concise, and that remain tailored to the nature of each target task.
Figure 3. An overview of the SEER framework, illustrating its self-enhancing procedure of generation, selection, and adaptive filtering.
4.1. Pre-Inference Data Generation
The first stage of SEER focuses on constructing a comprehensive corpus of CoT outputs to serve as the foundation for subsequent filtering and fine-tuning. In this step, the base language model is used to process each question in the training set, generating a complete solution that includes an explicit, step-by-step reasoning path.
A distinctive advantage of our approach is its low sensitivity to prompt engineering. While a prompt is required to elicit CoT reasoning, our subsequent adaptive filtering stage ensures that the method is not critically dependent on the choice of prompt format. This property allows us to adopt a pragmatic strategy: whenever an official, dataset-specific prompt template is available, we use it directly; otherwise, we revert to a robust general-purpose template. For example, for mathematical reasoning tasks, we utilize the system prompt from the study by Wei et al. (Wei et al., 2022):
[…]
To prevent incomplete reasoning caused by premature output truncation, we allocate a sufficiently high maximum token budget during generation. This ensures that the model’s natural CoT length distribution is faithfully captured. Accurately profiling this distribution is critical for calibrating the adaptive filter in later stages, as it allows us to set dataset-specific compression thresholds that minimize information loss.
4.2. Best-of-N Sampling for Data Refinement
To prevent the reasoning loops noted in Section 3, which can harm data quality, we adopt a BoN sampling strategy to build a high-quality training set. This strategy also helps obtain cleaner datasets with more concise CoT lengths. For each input question, the model generates $\mathcal{N}$ candidate responses, each augmented with a CoT. These candidates are then ranked and filtered according to a strict hierarchical set of criteria:
  1. (1)
    Correctness of the Final Answer: Only candidates producing the correct final answer are considered. All incorrect outputs are discarded.
  2. (2)
    Presence of Valid CoT: From the remaining candidates, we retain only those containing a non-empty reasoning path. This guarantees that each example is explainable and preserves its interpretability.
  3. (3)
    Conciseness of Reasoning: If multiple candidates satisfy the first two criteria, we select the one with the shortest CoT length. This prioritizes efficiency and prevents the inclusion of unnecessarily verbose reasoning.
This multi-stage filtering process ensures that the resulting dataset consists exclusively of reasoning traces that are (1) correct in outcome, (2) logically explicit, and (3) concise. By integrating BoN sampling into our pipeline, we are able to suppress reasoning loops, improve data quality, and provide the model with more effective fine-tuning signals.
4.3. Adaptive CoT Filtering
Even after BoN sampling, CoT responses can vary widely in length depending on the task, domain, and problem complexity. Excessively long CoTs waste computational resources, while overly aggressive compression risks discarding essential reasoning steps. To balance these factors, we introduce an adaptive CoT filtering mechanism that sets a dataset-specific maximum CoT length, denoted $\lambda_{c}$.
A fixed, manually chosen $\lambda_{c}$ is often suboptimal because optimal CoT lengths differ substantially across domains (e.g., mathematical proofs vs. code debugging). If $\lambda_{c}$ is too low, important intermediate reasoning may be lost, potentially reducing accuracy and even causing the model to overfit to overly compressed reasoning patterns. Conversely, a value set too high will dilute the benefits of compression, leading to minimal gains in efficiency.
Our adaptive filtering addresses this by first generating CoT responses for the entire training set during the pre-inference stage and then computing two statistical measures of their length distribution: the mean length $\bar{\lambda}$ and the median length $\tilde{\lambda}$. We define $\lambda_{ada}$ as:
$\lambda_{ada}=\frac{\bar{\lambda}+\tilde{\lambda}}{2}$
This formula anchors the threshold to the observed characteristics of each dataset, balancing robustness (median) with representativeness (mean).
While $\lambda_{c}$ provides a principled baseline, our design remains flexible. Users can manually adjust it to match specific efficiency or accuracy requirements. This tunable control allows practitioners to trade off between computational savings and reasoning completeness, enabling SEER to adapt seamlessly across diverse reasoning scenarios.
4.4. Overall Procedure
Input: Training set $\mathcal{D}=\{(x_{i},y_{i}^{\star})\}$, base model $\mathcal{M}$, candidates per input $\mathcal{N}$
Output: The model $\mathcal{M}_{\text{adapt}}$ fine-tuned on the filtered dataset $\widetilde{\mathcal{D}}$
Stage 1: Pre-inference generation;
foreach $(x_{i},y_{i}^{\star})\in\mathcal{D}$ do
  Generate $\mathcal{N}$ CoT-augmented candidates with $\mathcal{M}$ under a fixed prompt;
  $r_{ij}\leftarrow\text{final response generated for }x_{i}$;
  $\text{CoT}_{ij}\leftarrow\text{reasoning trace (Chain-of-Thought) leading to }r_{ij}$;
  $\mathcal{C}_{i}\leftarrow\{\,(r_{i1},\text{CoT}_{i1}),\ldots,(r_{i\mathcal{N}},\text{CoT}_{i\mathcal{N}})\,\}$;
Stage 2: Best-of-N (BoN) sampling;
foreach $(x_{i},y_{i}^{\star}),\mathcal{C}_{i}$ do
  $\mathcal{S}_{i}\leftarrow\{(r,\text{CoT})\in\mathcal{C}_{i}\mid\textsc{Correct}(r,y_{i}^{\star})\ \wedge\ |\text{CoT}|>0\}$ if $\mathcal{S}_{i}\neq\emptyset$ then
     $(r_{i}^{\star},\text{CoT}_{i}^{\star})\leftarrow\arg\min_{(r,\text{CoT})\in\mathcal{S}_{i}}|\text{CoT}|$ Add $(x_{i},r_{i}^{\star},\text{CoT}_{i}^{\star})$ to $\mathcal{D}^{\prime}$;
Stage 3: Adaptive CoT filtering;
Compute length set $L\leftarrow\{\,|\text{CoT}_{i}^{\star}|\mid(x_{i},r_{i}^{\star},\text{CoT}_{i}^{\star})\in\mathcal{D}^{\prime}\,\}$;
$\bar{\lambda}\leftarrow\text{mean}(L),\ \tilde{\lambda}\leftarrow\text{median}(L)$;  $\lambda_{c}\leftarrow(\bar{\lambda}+\tilde{\lambda})/2$;
$\widetilde{\mathcal{D}}\leftarrow\{(x_{i},r_{i}^{\star},\text{CoT}_{i}^{\star})\in\mathcal{D}^{\prime}\mid|\text{CoT}_{i}^{\star}|\leq\lambda_{c}\}$;
Stage 4: Fine-tuning;
Fine-tune model on $\widetilde{\mathcal{D}}$ with SFT to obtain $\mathcal{M}_{\text{adapt}}$;
return $\mathcal{M}_{\text{adapt}}$;
Algorithm 1 Self-Enhancing Training Process with BoN & Adaptive CoT Filtering
Algorithm 1 summarizes the end-to-end process, from pre-inference generation to BoN-based selection, adaptive filtering, and final fine-tuning on the curated dataset.
Through the combination of high-coverage pre-inference generation, strict BoN filtering, and adaptive CoT length control, SEER produces training data that is both high-quality and efficiency-oriented. This not only improves the reasoning efficiency of the fine-tuned model but also mitigates failure modes such as infinite reasoning loops and truncation.
5. Experimental Setup
5.1. Datasets
To thoroughly evaluate the performance and efficacy of our method, we conduct experiments across three distinct software engineering tasks and one mathematical reasoning task. These tasks were selected to cover a broad spectrum of problem types, input–output modalities, and reasoning requirements. By incorporating both code-related and mathematical reasoning benchmarks, we aim to investigate not only the in-domain effectiveness of our method but also its adaptability to different reasoning paradigms and task formats.
  • MathQA-Python (Austin et al., 2021) is a dataset designed for generating Python programs that solve mathematically formulated problems. Each problem is presented in natural language and requires the model to produce an executable Python solution. The challenges in this dataset include parsing problem statements, performing intermediate symbolic reasoning, and implementing the reasoning steps as correct Python code. This dataset serves as a baseline for evaluating code generation in mathematically oriented problem-solving contexts.
  • CodeXGLUE-Defect-Detection (Zhou et al., 2019) is a binary classification dataset that assesses a model’s ability to identify whether a given piece of source code contains security vulnerabilities or functional defects. The data originates from real-world code bases, and the detection task often requires understanding complex control flows, data dependencies, and security implications. This dataset evaluates the model’s capability in static code analysis and secure software engineering scenarios, where concise yet accurate reasoning is critical to avoid excessive inference overhead.
  • Code-Search-WebQuery (Lu et al., 2021; Husain et al., 2019) is used to test semantic code retrieval capabilities. In this task, the model is provided with a natural language query and a code snippet, and must determine whether the code correctly satisfies the described requirement. This problem is framed as binary classification, and it challenges the model to bridge the semantic gap between natural language descriptions and programming logic. The task reflects realistic software development workflows, such as searching for relevant functions in large code bases.
  • GSM-8K (Cobbe et al., 2021) is a large-scale benchmark for multi-step mathematical reasoning. Each question requires systematic decomposition into intermediate steps before arriving at the final answer. Unlike MathQA-Python, the output here is not executable code but a final numeric answer, making it an ideal testbed for evaluating whether our method can compress reasoning chains without degrading accuracy in purely symbolic reasoning scenarios. It also provides an out-of-domain evaluation for models fine-tuned on code-related tasks.
Additionally, in RQ3 we include the HumanEval dataset introduced in Section 3 to evaluate the generalizability of our method. By selecting this combination of datasets, our evaluation covers tasks ranging from code synthesis, defect detection, and semantic code retrieval to mathematical reasoning.
5.2. Implementation Details
All experiments are conducted on a server equipped with four NVIDIA RTX A800 GPUs (80 GB each) and 128 Intel Xeon CPUs, running Ubuntu 22.04 LTS. We fine-tune DeepSeek-R1-Distill-Qwen-7B model (DeepSeek-AI, 2025b) under a full-parameter supervised fine-tuning (SFT) setting, using the LLaMAFactory framework (Zheng et al., 2024). Training is performed for three epochs with an effective batch size of eight, gradient checkpointing enabled, and the AdamW optimizer. The learning rate is fixed at $1\times 10^{-5}$ with a cosine decay schedule and 10% warm-up, and gradient clipping (norm 1.0) is applied for stability. During fine-tuning, inputs are tokenized with a maximum sequence length of 8,192 tokens, truncating from the input side if necessary to preserve the output reasoning chain. Although we adopt full-parameter SFT by default, the SEER framework is also compatible with parameter-efficient fine-tuning methods such as LoRA and QLoRA, which we discuss in the Discussion section.
5.3. Baselines
In our work, we consider the following baselines.
  • Self-Training: Directly using self-inference data to train the model.
  • TokenSkip (Xia et al., 2025): Fine-tuning LLMs by selectively skipping less important tokens while retaining key tokens, enabling the generation of compressed CoTs with adjustable ratios.
  • Naive BoN (Munkhbat et al., 2025b): Use only the Best-of-N method to process the pre-inference data and then train the model.
In order to ensure the fairness of the experiment, we have implemented several modifications to the TokenSkip source code. These changes include: 1) the incorporation of the “deepseek” prompt template, as this template is employed by the chosen model; and 2) the use of the vllm library in place of the transformers library to achieve more efficient inference and to mitigate any potential precision loss that could lead to unfair comparisons.
5.4. Evaluation Metrics
We measure the following metrics:
[…]
Accuracy: We use the same pass@1, which is mentioned in Section 3 as the primary metric to evaluate correctness. For programming tasks (e.g., MathQA-Python), a prediction is considered correct if the generated code passes all test cases. For mathematical reasoning tasks (e.g., GSM-8K), we adopt exact matching criterion, where the model’s final answer must exactly match the ground truth.
[…]
Compression Rate: We define the Compression Rate $\rho$ to quantify the conciseness of the generated CoT. Specifically, $\rho$ measures the relative reduction in the number of tokens compared to a reference CoT. It is computed as: $\rho=(\lambda-{\lambda^{\prime}})/{\lambda}$
In this formula, $\lambda$ denotes the token count of the reference (original) CoT, and $\lambda^{\prime}$ represents the token count of the CoT generated by the evaluated model or baseline. A higher value of $\rho$ indicates a greater degree of compression, corresponding to a more concise reasoning path. This metric is particularly important for evaluating methods such as TokenSkip and Naive BoN, which are explicitly designed to produce shorter CoTs without compromising correctness.
6. Experiment
In this section, we conduct experiments to answer the following research questions.
6.1. Research Questions
In order to evaluate our work, we address the following research questions:
  • RQ1: How does the CoT compression performance of SEER compare with baselines? We evaluate SEER against baseline methods on multiple datasets.
  • RQ2: What is the contribution of each component of SEER? We conduct ablation studies by removing or varying parts of SEER, such as the CoT length filter $\lambda_{c}$ and the BoN parameter $\mathcal{N}$, to examine their impact.
  • RQ3: How well does SEER generalize across domains? We examine whether SEER remains effective when the fine-tuned data and test data come from different domains.
  • RQ4: Can SEER mitigate reasoning loops? We test whether SEER helps LLMs generate valid answers without falling into infinite reasoning loops.
6.2. RQ1: The Performance of SEER
In this subsection, we assess the performance of SEER in comparison with baseline methods. For the TokenSkip experiment, we set $\gamma=0.5$, meaning that we append the sequence “[EOS]0.5[EOS][EOS]0.5[EOS]” to the end of the input prompt. Similarly, for Naive BoN, we set $N=3$, which is consistent with the parameters used in evaluating our method. Table 2 reports task pass@1 accuracy ($\mathcal{A}$), average CoT length ($\lambda_{\text{avg}}$), and compression rate ($\rho$) under the 8K token constraint.
Methods MathQA-Python Defect-Detection Code-Search GSM-8K $\rho_{avg}(\%)$
$\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$
Base 63.7 1,456 - 44.7 1,836 - 72.4 472 - 90.1 1,309 - -
Self-Trained 74.6 1,177 19.2 49.3 1,055 42.5 72.5 468 0.9 89.2 792 39.5 25.5
Naive BoN (N=3) 73.6 1,191 18.2 49.8 1,042 43.3 74.9 478 -1.3 89.8 845 35.5 23.9
TokenSkip ($\gamma$=0.5) 1.6 1,196 17.9 36.2 2,421 -31.9 66.5 465 1.5 37.9 1,238 5.4 -1.8
TokenSkip ($\gamma$=0.8) 0.9 1,109 23.8 47.2 1,004 45.3 76.4 341 27.8 87.3 1,197 8.6 26.4
SEER 74.9 877 39.8 50.5 785 57.2 77.3 362 23.3 90.5 677 48.3 42.1
    The Base model refers to “DeepSeek-R1-Distill-Qwen-7B”. The same applies to the other tables;
    $\mathcal{A}(\%)$: Pass@1; $\lambda_{avg}$: Average output length (tokens);
    $\rho(\%)$: Compression rate compared to the Base model; $\rho_{avg}(\%)$: Average compression rate across all datasets.
Table 2. The CoT Compression Performance Over Different Datasets (8K Token Limit)
Across all three software engineering tasks, SEER shows strong performance in both accuracy and CoT compression. On MathQA-Python, which requires code generation, it achieves 74.9% pass@1 with a 39.8% reduction in reasoning length. On Defect-Detection, it obtains the best pass@1 (50.5%) and the highest compression rate (57.2%). On Code-Search, it maintains a high pass@1 of 77.3% while compressing CoTs by 22.9%. Beyond software engineering, SEER also performs well on GSM-8K, a math reasoning dataset, reaching 90.5% pass@1 with a 48.3% reduction in CoT length. The accuracy gains are largely due to the mitigation of truncation and excessive reasoning, which often undermine model reliability. These results demonstrate that SEER consistently balances brevity and accuracy across diverse tasks.
[…]
Compared to TokenSkip, both Self-Trained and Naive BoN achieve more stable outcomes but still fall short of SEER. On Defect-Detection, their pass@1 remains below 50%, with compression rates under 30%. While they provide modest improvements over the base model, they cannot match SEER’s combination of high accuracy and substantial compression. This suggests that generic fine-tuning or simple BoN sampling is insufficient for the challenging trade-off between performance and efficiency.
[…]
Answer to RQ1: Compared with all baselines, SEER consistently achieves high accuracy and efficient CoT compression on both software engineering tasks and the math reasoning task. On average, SEER reduces CoT length by 42.1%, representing a 16.9% improvement over the best existing baseline, while preserving or even improving accuracy across diverse tasks.
6.3. RQ2: The Impact of Each Components
To investigate the contribution of each component, we conduct ablation studies on the MathQA-Python dataset by varying the BoN sampling size $\mathcal{N}$ and the CoT length filter threshold $\lambda_{c}$. The results are summarized in Table 3.
Table 3. The Ablation Results of SEER on MathQA-Python Dataset
Methods Settings Pass@1 $\mathcal{A}$ (%) Avg CoT Len $\lambda_{avg}$ (Tokens) Compression Rate $\rho$ (%)
Base - 63.67 1,456 -
Without Filter $\mathcal{N}=1$ 73.49 1,191 18.20
$\mathcal{N}=3$ 74.45 1,177 19.16
$\mathcal{N}=5$ 74.40 1,163 20.12
Without BoN $\lambda_{c}=2\times\lambda_{ada}$ 74.40 1,071 26.44
$\lambda_{c}=1.5\times\lambda_{ada}$ 74.19 918 36.95
$\lambda_{c}=\lambda_{ada}$ 73.50 878 39.70
$\lambda_{c}=0.5\times\lambda_{ada}$ 72.76 836 42.58
SEER
$\mathcal{N}=3$ $\lambda_{c}=\lambda_{ada}$
74.88 877 39.77
When only BoN sampling is applied without the filter, increasing $\mathcal{N}$ from 1 to 3 improves accuracy from 73.49% to 74.45%, while the average CoT length decreases slightly from 1,191 to 1,177 tokens. However, further increasing $\mathcal{N}$ to 5 does not provide additional benefit and even slightly reduces accuracy (74.40%), indicating that the advantage of BoN sampling quickly saturates. In this setting, the compression rate remains relatively modest, ranging from 18.20% to 20.12%.
In contrast, applying only the adaptive filter leads to a substantial increase in compression. For instance, when $\lambda_{c}=2\times\lambda_{ada}$, the compression rate reaches 26.44%, and as the threshold becomes stricter ($\lambda_{c}=0.5\times\lambda_{ada}$), compression improves further to 42.58%. Nevertheless, overly aggressive filtering reduces accuracy, dropping from 74.40% to 72.76%. These results show that while the filter is more effective than BoN in reducing the redundancy in CoT, it risks discarding essential reasoning steps when applied too strongly.
To balance accuracy and compression rate, we set $\mathcal{N}=3$ and $\lambda_{c}=\lambda_{ada}$ in our method. When both components are combined in SEER with these settings, the method achieves the best trade-off, with the highest accuracy of 74.88% and a competitive compression rate of 39.77%. This demonstrates that the two modules are complementary: BoN sampling enhances robustness, while the adaptive filter enforces compactness, and together they enable SEER to deliver concise yet accurate reasoning.
[…]
Answer to RQ2: The ablation results reveal that BoN sampling alone offers only limited gains, while the adaptive filter strongly enhances compression but may harm accuracy if too strict. Their combination achieves the best trade-off, yielding both higher accuracy and effective compression, highlighting the complementary roles of the two components.
6.4. RQ3: Generalizability of SEER
In this section, we evaluate the generalizability of SEER by fine-tuning it on one dataset and testing on the remaining three datasets. The results are presented in Table 4.
Table 4. The Generalizability Evaluation Result of SEER Across All Tasks
MathQA-Python Defect-Detection Code-Search GSM-8K
Models $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$ $\mathcal{A}(\%)$ $\lambda_{avg}$ $\rho(\%)$
Base 63.7 1456 - 44.7 1,836 - 72.4 472 - 90.1 1,309 -
MathQA-Py Fine-tune 74.9 877 39.8 47.1 1131 38.4 77.3 395 16.3 87.2 748 42.9
Defect-Detection Fine-tune 64.2 1279 12.2 50.5 785 57.2 72.2 339 28.2 90.4 1128 13.8
Code-Search Fine-tune 64.1 1104 24.2 51.4 672 63.4 77.3 362 23.3 89.8 853 34.8
GSM-8K Fine-tune 62.5 1138 21.8 48.5 1483 19.2 79.6 435 7.8 90.5 677 48.3
Overall, the findings demonstrate that SEER exhibits strong cross-task generalization. Regardless of which dataset is used for fine-tuning, the resulting models generally achieve higher accuracy compared with the base model. Only a few cases show a slight decrease, such as fine-tuning on GSM-8K and then evaluating on MathQA-Python (62.5% vs. 63.7%). In all cases, the compression rate remains consistently above zero. For example, fine-tuning on MathQA-Python compresses average reasoning length on GSM-8K from 1,309 tokens to 748 tokens (42.9% compression), while still preserving high accuracy (87.2%). This indicates that SEER is able to reduce reasoning length while maintaining competitive accuracy even on out-of-domain tasks.
Interestingly, in some settings, the fine-tuned models even surpass those trained directly on their target dataset. For instance, the model fine-tuned on Code-Search achieves 51.4% accuracy on Defect-Detection, which is higher than the model fine-tuned on Defect-Detection (50.5%). These results suggest that SEER captures reusable reasoning patterns that extend beyond the dataset it was trained on, rather than overfitting to a single task.
Figure 4. Performance of SEER fine-tuned on different datasets when evaluated on HumanEval. Fine-tuning on coding-related datasets such as MathQA-Python and Code-Search yields the best generalization with shorter CoT lengths, while non-coding datasets like GSM-8K still provide moderate improvements over the base model.
To further assess its generalizability, we evaluate the four fine-tuned models on the HumanEval dataset, as shown in Figure 4. The results reveal two key insights. First, all fine-tuned models achieve noticeable improvements in both accuracy and CoT length over the base model, confirming that the knowledge learned from different tasks generalizes to program synthesis. Second, the reduction in reasoning length remains consistent, suggesting that SEER’s compression mechanism is stable even when applied to tasks outside the training domain. This demonstrates that the framework does not merely adapt to a specific dataset but instead learns reasoning strategies that generalize across diverse software engineering and mathematical tasks.
[…]
Answer to RQ3: SEER demonstrates strong cross-task generalization. Even when evaluated on out-of-domain datasets, it consistently improves accuracy while reducing CoT length. Fine-tuning on coding datasets further generalizes well to HumanEval, confirming that SEER learns reusable reasoning strategies beyond its training domain.
6.5. RQ4: The Ability of Reducing Reasoning Loops
Infinite reasoning loops occur when a model repeatedly generates identical reasoning segments until the output reaches the maximum context length and is truncated. This phenomenon wastes computational resources, inflates latency, and often produces incomplete or invalid final answers. Because such loops can arise unpredictably across models of different sizes, they represent a critical robustness concern. In this section, we investigate whether our approach can mitigate this problem.
For each dataset, we adapt our method to fine-tune a model specific to that dataset and evaluate it against the base model. The results in Table 5 show that SEER substantially reduces reasoning loops in all cases. On MathQA-Python, loop occurrences drop from 85 to 23 (72.9% reduction). On Code-Search, the count falls from 15 to 1 (93.3% reduction), effectively eliminating loops. The largest improvement appears in Defect-Detection, where loops decrease from 222 to just 7 (96.8% reduction). Even on GSM8K, which generally contains shorter reasoning chains, loops drop from 54 to 6 (88.9% reduction).
Table 5. The Performance of SEER in Reducing Truncation and Reasoning Loops
MathQA-Python Code-Search Defect-Detection GSM8K
Truncation Loops Truncation Loops Truncation Loops Truncation Loops
Base 110 85 18 15 235 222 59 54
SEER 30 23 1 1 7 7 8 6
During our empirical analysis in Section 3, we find that most truncations in the base model are directly caused by such loops. This behavior wastes much of the available context budget and prevents the model from completing its reasoning. Our adaptive filtering strategy compresses CoT length, which mitigates loop formation and reduces truncation, as shorter reasoning provides fewer chances for repetitive output. This dual effect, loop suppression and truncation reduction, further enhances the robustness and practicality of our method in real-world deployments.
Overall, these results confirm that SEER offers a robust defense against infinite reasoning loops, complementing its primary goal of efficient CoT compression and providing benefits for latency, reliability, and output quality.
[…]
Answer to RQ4: SEER effectively suppresses infinite reasoning loops across all datasets, with reductions ranging from 72.9% on MathQA-Python to 96.8% on Defect-Detection. Because most truncations in the base model stem from such loops, the adaptive filtering strategy not only prevents pathological repetition but also lowers truncation frequency.
7. Discussion
7.1. The Impact of Training Approaches
An important aspect of our framework is its adaptability to different fine-tuning paradigms. Beyond full-parameter supervised fine-tuning (SFT), we also explored parameter-efficient alternatives such as LoRA. The comparison reveals two insights. First, while SFT unsurprisingly achieves the strongest overall performance, LoRA retains most of the benefits of our method, delivering substantial accuracy gains and compression improvements over the base model. This indicates that the effectiveness of adaptive CoT filtering does not depend on updating the entire parameter set.
Table 6. The Performance Comparison Between Using SFT and PEFT Fine-tuning Method
Pass@1 $\mathcal{A}$ (%) Avg CoT Len $\lambda_{avg}$ (Tokens) Compression Rate $\rho$ (%)
Base 63.67 1,456 -
PEFT (LoRA) 67.71 949 34.8
SFT 74.88 877 39.8
Second, the LoRA-based approach demonstrates that our approach remains effective in environments with limited resources. By reducing computational and storage costs, LoRA enables deployment in scenarios where full SFT would be prohibitively expensive. Although the absolute gains are somewhat smaller than those under SFT, the efficiency–effectiveness trade-off makes PEFT methods an attractive option for real-world applications.
Taken together, these observations indicate that our framework is effective under ideal training conditions and remains flexible in limited-resource settings. This adaptability broadens its applicability, making it feasible to integrate SEER into a wider range of deployment scenarios.
7.2. Threats to Validity
7.2.1. Internal Validity
Internal validity concerns whether the observed improvements of SEER truly arise from our proposed framework rather than confounding factors. One potential threat is the reliance on a single model family, namely the DeepSeek-R1-Distill-Qwen-7B. Although we carefully fine-tuned and evaluated under consistent settings to reduce bias, the conclusions may partially reflect properties of this architecture rather than general effectiveness. To mitigate this risk, we performed ablation studies (varying BoN sampling and adaptive filtering thresholds) to confirm that the observed gains stem from the specific design choices of SEER rather than incidental artifacts. Future work will extend the evaluation to other model families, including implicit CoT models that do not explicitly output reasoning traces. Moreover, testing on closed-source models with advanced reasoning capabilities (e.g., OpenAI O1, Gemini, Claude’s Thinking series) will help validate whether the improvements are architecture-independent.
7.2.2. External Validity
External validity concerns the generalizability of our findings beyond the datasets and settings considered in this work. Our experiments primarily focus on software engineering tasks, as they represent one of the most critical application domains for reasoning models. To mitigate the risk of over-specialization, we additionally included a mathematical reasoning dataset (GSM-8K), which provides a different reasoning style and serves as an initial step toward broader validation. Nevertheless, important domains such as commonsense reasoning, legal QA, and scientific problem solving remain unexplored.
Moreover, all evaluations were conducted with fixed prompts and prompting strategies. While our method is designed to be prompt-agnostic, alternative instruction formats, few-shot configurations, or domain-specific prompts could influence the observed trade-off between compression and accuracy. Finally, although the datasets used contain tasks of varying difficulty, we have not systematically examined whether compression benefits differ across problem complexity levels.
To further strengthen external validity, we plan to expand our evaluation to additional reasoning domains and more diverse datasets, as well as investigate robustness under alternative prompting strategies. Future work will also include stratified analyses by task difficulty to better understand whether SEER is equally effective on both simple and challenging instances.
8. Related Works
Since the concept of CoT was first introduced (Wei et al., 2022), research on controlling and optimizing CoT length has drawn increasing attention. Numerous studies have shown that extending intermediate reasoning steps within CoT prompting can significantly enhance the capabilities of LLMs (Fu et al., 2022; Merrill and Sabharwal, 2023; Jin et al., 2024). However, other work (Wu et al., 2025) indicates that performance does not increase indefinitely; instead, it follows a rise-and-fall pattern as CoT length grows. This suggests that optimally controlling CoT length is a key challenge for balancing reasoning quality and efficiency.
$\blacksquare$ Prompt-Guided Efficient Reasoning. Prompt-based approaches improve reasoning efficiency by explicitly guiding the model to be concise. CCoT (Cheng and Van Durme, 2024) uses a concise CoT template to encourage brevity through prompt wording. Token-Budget (Han et al., 2024) proposes a token-budget-aware reasoning framework that dynamically adjusts reasoning length based on estimated complexity. CoD (Xu et al., 2025a) limits reasoning to within five words via a “Chain-of-Draft” template. Token Complexity (Lee et al., 2025) systematically analyzes prompt compression strategies, highlighting that many still operate far from theoretical efficiency limits.
These methods can reduce verbosity without modifying model parameters, but they are inherently sensitive to prompt design and task distribution. In contrast, our method removes this dependency: by learning from filtered, concise, and correct CoT examples, the model internalizes concise reasoning patterns that generalize across prompts and domains.
$\blacksquare$ Model-Optimization Based Efficient Reasoning. Other approaches modify the model or its training data to achieve concise reasoning. C3oT (Kang et al., 2025) trains models jointly on long and short CoTs to learn mappings between them, using conditioned inference to produce short CoTs. TokenSkip (Xia et al., 2025) enables skipping less important tokens during generation, allowing controllable compression but sometimes causing instability (e.g., reasoning loops in our observations). LightThinker (Zhang et al., 2025) compresses verbose thoughts into compact internal representations, discarding the original CoT from the context. CODI (Shen et al., 2025) distills CoT into a continuous space, jointly learning explicit and implicit reasoning traces.
While these methods reduce reasoning length, many require additional compression modules, complex training objectives, or implicit representations that limit interpretability. Our method is simpler: it uses the model’s own generated data, combined with a lightweight Best-of-N sampling and adaptive filtering, to directly fine-tune for concise yet explicit reasoning. This avoids extra components and preserves the transparency of reasoning steps.
$\blacksquare$ Chain-of-Thought Effectiveness in Software Engineering. CoT has shown strong benefits in software engineering tasks. SCoT (Li et al., 2025) was among the first to apply structured CoT to code generation. The emergence of Deepseek-R1 (DeepSeek-AI, 2025b) further highlighted the value of streamlined CoT for software engineering tasks. Liu et al. (Liu et al., 2025) revisited CoT training paradigms, showing that some seemingly effective methods do not transfer well to real-world coding tasks. UnCert-CoT (Zhu et al., 2025) uses uncertainty estimation to decide when to trigger CoT, avoiding unnecessary reasoning. CoD (Yang, 2025) adapts the “Chain-of-Draft” approach to software engineering tasks, greatly reducing token usage.
However, existing SE-focused methods often target specific task scenarios or rely on inference-time control, which may not eliminate pathological failure modes, such as infinite reasoning loops. Our method differs by explicitly filtering out redundant or looping reasoning during training, allowing the model to internalize concise, stable reasoning behaviors. This makes it robust not only in software engineering tasks but also in mathematical reasoning, while retaining interpretability and avoiding the need for task-specific prompt engineering.
9. Conclusion
Our work systematically examined the inefficiency of excessively long CoT reasoning, identifying that long CoT can lead to truncation and overthinking, resulting in decreased accuracy. We then present SEER, a self-enhancing framework for adaptive CoT compression. Through extensive experiments on software engineering and mathematical reasoning tasks, SEER reduces CoT length by an average of 42.1% while maintaining or even improving accuracy. Our ablation results confirm the complementary roles of BoN sampling and adaptive filtering, and cross-domain evaluations demonstrate SEER’s strong generalizability. SEER also effectively mitigates truncation and suppresses infinite reasoning loops, addressing key efficiency and robustness issues. These findings highlight SEER as a practical approach for making CoT-enhanced LLMs more efficient and reliable. Future work will extend SEER to broader reasoning domains, apply it to diverse model architectures, and explore integration into latency-sensitive real-world systems.